Sphere A has a charge of +10.0 C.It is brought in contact with a neutral sphere B and then separated.Find the final charges on spheres A and B if the spheres have equal radius and if the radius of sphere B is twice the radius of sphere A


1
Expert's answer
2022-02-17T04:36:25-0500

Given:

"q_1=+10.0\:\rm C"

"q_2=0\:\rm C"


The law of conservation of a charge says

(a)

"q_1+q_2=q_1'+q_2'"

"q_1=q+q=2q"

"q_1'=q_2'=q=q_1/2=(+10.0\:\rm C)/2=+5.0\: C"

(b)

"q_1+q_2=q_1'+q_2'"

"q_1=q+2q=3q"

"q_1'=q=q_1/3=(+10.0\:\rm C)/3=+3.33\: C"

"q_2'=2q=2q_1/3=2(+10.0\:\rm C)/3=+6.67\: C"